OpenStudy (anonymous):

CAN SOMEONE CHECK IF MY ANSWERS CORRECT The ratio of men to women working for a company is 2 to 3. If there are 130 employees total, how many women work for the company?

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } = \frac{ 130 }{ x }\] \[2x=130 \times 3\] \[x = \frac{ 130 \times 3 }{ 2}\] \[x = 195\]

Vocaloid (vocaloid):

not quite.

Vocaloid (vocaloid):

hold on a sec

OpenStudy (anonymous):

can you help me fix it ? what did i do wrong ?

Vocaloid (vocaloid):

there are 2 men for every 3 women, so would you agree that out of every 5 employees, 2 are men and 3 are women?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

would you agree that 2/5 of the employees are men and 3/5 of the employees are women, then?

OpenStudy (anonymous):

yes i would

Vocaloid (vocaloid):

ok, so we know that 3/5 of the employees are women there are 130 total employees can you find how many women there are?

OpenStudy (anonymous):

is 78 correct ?

Vocaloid (vocaloid):

yup! you got it!

OpenStudy (anonymous):

thank you !